OpenStudy (bloomlocke367):

Okay, I just want to make sure I'm doing this correctly. Would I put 80 over the total (400) and multiply that by 2?

OpenStudy (bloomlocke367):

@rational

OpenStudy (anonymous):

Yes, you are doing this correctly

OpenStudy (bloomlocke367):

@rational

OpenStudy (bloomlocke367):

oh, okay. thanks XD

OpenStudy (bloomlocke367):

80/400=1/5. 1/5*2=2/5

OpenStudy (bloomlocke367):

that's not an option... what did I do wrong?

OpenStudy (rational):

the probability is just about drawing a blazers student in the "second try"

OpenStudy (bloomlocke367):

ohhh, so it's just 1/5?

OpenStudy (rational):

in the second try also 80 blazers students and 400 total students so the probability is simply 80/400

OpenStudy (rational):

Yes!

OpenStudy (bloomlocke367):

wait, but it asked the prob the principal will draw ANOTHER student...

OpenStudy (bloomlocke367):

in that, wouldn't there be the possibility that the same student would be drawn?

OpenStudy (rational):

yes "another blazers student" he can be same as the previous one

OpenStudy (bloomlocke367):

ohhhh okay.

OpenStudy (bloomlocke367):

thank you c: I really appreciate it.

OpenStudy (rational):

np :) that "ANOTHER" word got me too